OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence.a16 = 21, a17 = -1

OpenStudy (anonymous):

@dumbcow @nincompoop @IMStuck @zepdrix @waterineyes

OpenStudy (anonymous):

-(-8-6t)

OpenStudy (anonymous):

@WilliamZhang can you help?

OpenStudy (anonymous):

A16=21, A17=-1, and it is an arithmetic sequence. So d= A17-A16=-1-21=-22 A1=A16-15d=21+15*22=251 So An=A1+(n-1)d=251-22(n-1)=273-22n